OpenStudy (anonymous):

25, 34, 46, 37, 33, 42, 40, 37, 34, 49, 73, 46, 45, 45, 5 The first quartile of this distribution, 34, is also called the 25th percentile, indicating that approximately 25% of the observations fall below this value. The third quartile of this distribution, 46, is also called the 75th percentile, indicating that approximately 75% of the observations fall above this value. In this distribution, what percent of the observations actually fall outside of the IQR? That is, what percent of the observations fall below 34

OpenStudy (paxpolaris):

re-writing in order: 5, 25, 33, 34, 34, 37, 37, 40, 42, 45, 45, 46, 46, 49, 73 15 numbers in total what percent of the observations actually fall outside of the IQR? That is, what percent of the observations fall below 34 ... or above 46? 5,25,33, 49 and 73 so 5 out of 15 or 1/3.

OpenStudy (paxpolaris):

=33.3%

OpenStudy (anonymous):

My math teach said : 20%

OpenStudy (paxpolaris):

outside the IQR means: below 34 OR above 46 if it was specifically asking for JUST below 34: then there are just 3 numbers: 5, 25 and 33. 3/15 = 1/5 = 20% outside the IQR does not just mean below 34.

OpenStudy (anonymous):

ohhhh okkay, thankyou! :)